Altawfik Corporation is estimating the cost function for total cost of production of product A using the two points method.  The data collected for the past year is as following:

                                       Number of units        Total

                Quarter            produced                   Costs

                      1                     4,000                $   1,000

                      2                     5,400                     1,280

                      3                     7,000                     1,600

                      4                     9,000                     2,000

            Using the two points methods, choose two points with which you estimate the cost function.
Answer #1

It doesn't matter which two points you choose, you will always get the same function.

Now, cost function is C(x)= V(x)+FC,

where V=Variable cost per unit

x=units

FC= Fixed costs

Now variable cost per unit = ACost/ A Unit

Taking quarter 1 and 2, Variable cost/unit= 280/1400= $0.2 (Take any quarter and you will get the same answer).

now C(x)=V(x)+FC

Taking quarter 1, 1000=4000*0.2+FC

FC= $200.

So, your cost function is -> C(x)=0.2x+200.